OpenStudy (anonymous):

if y' = 2x-4, which of the following is an equation for this curve if it passes through point (2,2)?

OpenStudy (anonymous):

a) y = 2x -2 b) y = x^2 -2 c) y = x^2-4x d) y = x^2-4x+6 e) y = 2x^2-4x+2

OpenStudy (anonymous):

Have you done integration in your class yet?

OpenStudy (anonymous):

nope, we only learned power rule

OpenStudy (anonymous):

Well, makes things more difficult, but okay. So, would you know how to differentiate all your options, a through e?

OpenStudy (anonymous):

yes. a = 2 b = 2x c = 2x - 4 d = 2x - 4 e = 4x^3-4

OpenStudy (anonymous):

Well, the last one is 4x-4, but Im sure you knew that :P So clearly only c and d could work. In order to see which one, you would need to plug in x = 2 and see if you get y = 2 c: (2)^2 - 4(2) = ? d: (2)^2 - 4(2) + 6 = ? So which one of those gives you a y value of 2? Because the coordinate point basically says that when you plug in 2, you get an answer of 2

OpenStudy (anonymous):

oooh so the answer is d, thanks!!!

OpenStudy (anonymous):

Yep, np :)